# Product of Experts for Visual Generation

### v1

@@ D.1 Flow Models
We require that $\xi$ is monotonic and $\xi(T)=0,\xi(1)=1$. By construction, $p^{(i)}_{T}(x)=\mathcal{N}(0,I)$ is easy to sample from, and $p^{(i)}_{1}(x)=p_{\text{data}}^{(i)}(x)$ encapsulates the prior data distribution of the expert model, fulling our motivation in Section 3.1. For all experiments, we define $\xi$ to be the Euler discretization esser2024scaling ; euler1792institutiones , with additional alignment to the DDPM scheduler ho2020denoising for experiments in Section 4.4.
@@
$s_{t}^{(i)}(x)$ $=\frac{\dot{\alpha}_{t}}{\alpha_{t}}x-\frac{\dot{\sigma_{t}}\sigma_{t}\alpha_{t}-\dot{\alpha}_{t}\sigma^{2}_{t}}{\alpha_{t}}v_{t}^{(i)}(x),$
@@
where $\lambda^{(i)}(x)$ is defined in Section 4.1, with all entries of $\sum_{i}\lambda^{(i)}(x)$ to be $1$.
D.2 Approximation for Conditional Sampling
We explain Eq. 7 below. Let $x_{i}$ be any expert, we then approximate the distribution $p(x_{i}\mid x_{\text{pa}(i)})$ as
@@ D.2 Approximation for Conditional Sampling
Appendix E Extended Experimental Results
E.1 Qualitative Video Results
Please view the “index.html” in the supplementary material folder using a web browser.
E.2 Intermediate Sample Visualizations
@@
The parallel sampling process is visualized in Fig. 8. For each scene from Section 4.4, we use an LLM to convert a text prompt into a layout of bounding boxes following the procedure described in Section F.4. The input text prompts and resulting layouts and regional prompts are visualized in Fig. 6. Then, we follow Algorithm 1 and initialize $L=4$ particles and annealed schedule length $T=28$, which are reweighed at three key timesteps $t=21,19,1$ using the discriminative expert, VQAScore lin2024evaluating , which measures image-text alignment with a probability score. At $t=21,19$, we binarize the scores with the threshold set to be the median score of the current 4 particles, discard the two below the threshold, and duplicate the remaining two high-score particles two times to maintain 4 particles in each round. At $t=1$, we keep the highest-scoring particle as the final output sample.
@@
Fig. 8 shows end-point prediction, $\hat{x}^{(l)}=\frac{\sigma_{t}}{\dot{\alpha}_{t}\sigma_{t}-\dot{\sigma}_{t}\alpha_{t}}v_{t}(x^{(l)})-\frac{\dot{\sigma}_{t}}{\dot{\alpha}_{t}\sigma_{t}-\dot{\sigma}_{t}\alpha_{t}}x^{(l)}$, to visualize the intermediate samples from the annealing process from $t=28$ to $t=1$. VQAScore measurements (higher means more aligned) are overlayed on the top-left corner for each sample. In the proposed framework, integrating multiple generative experts allows for composing local, simpler regional prompts, construction of the annealed distributions allows for iterative refinement of samples, and parallel sampling further integrates the knowledge from the discriminative expert and improves output accuracy, as evidenced in the comparisons from Fig. 6.
Appendix F Implementation Details
F.1 Automatic Evaluation with VLMs
@@
We follow the protocol proposed in PhysGen3D chen2025physgen3d to run automatic evaluation for Sections 4.4 and 4.3 using GPT-4o hurst2024gpt4o . For each scene, all methods’ output images, or uniformly sampled 10 frames when outputs are videos, are sent to GPT-4o together with template prompts as specified in PhysGen3D chen2025physgen3d . GPT-4o is asked to give scores from 00 to $1$ for each video on all metrics. The scores for the same metric are averaged across all scenes.
F.2 Graphics-Engine-Instructed Image Editing
@@
In Section 4.2, the product distribution is defined as $p(x)\propto p^{\text{depth2image}}(x\mid c^{\text{text}},c^{\text{depth}})p^{\text{imagefill}}(x\mid c^{\text{text}},c^{\text{image}},c^{\text{mask}})$. We use $T=28,L=1$, with 1 MCMC sampling step and 2 conditional sampling updates (Eq. 7) per iteration, and spatial resolution $960\times 1664$.
F.3 Physical-Simulator-Instructed Video Generation
@@
In Section 4.3, the product distribution is defined as $p(x)\propto p^{\text{depth2video}}(x\mid c^{\text{text}},c^{\text{depth}})p^{\text{image2video}}(x\mid c^{\text{text}},c^{\text{image}})$. We use $T=30,L=1$ with 1 MCMC sampling step per iteration. We use no conditional sampling updates for this task due to GPU memory constraint. Each generated video has $81$ frames and spatial resolution $480\times 832$.
F.4 Text-to-Image Generation with Regional Composition
Text-to-Layout.
We use an LLM (GPT-o1-mini) to convert an input text prompt into a global caption, a list of region captions, and a layout specified by bounding boxes through the following steps.
We query LLM to separate the input text prompt into a list of main regions, each associated with a region caption. Then we pass the original prompt along with the separated list of region captions and ask it to assign a spatial box to each region, specified as a tuple of percentages: (start_x, end_x, start_y, end_y), where each value is in the range $[0,100]$. We instruct that these boxes are intended to form a plausible image layout that must satisfy all spatial constraints described in the prompt. We instruct to make each box as large as possible, subject to the following constraints: the width and height must each be at least 40; the width must be between 0.6×0.6$\times$0.6 × and 1.5×1.5$\times$1.5 × the height; neighboring boxes should overlap unless the prompt explicitly specifies that objects should be separated; however, the overlap between any two boxes must not exceed $50\%$.
Layout-to-Image Generation.
We use regional generative experts and discriminative experts for layout-to-image generation, or $p(x)\propto\left(\prod_{i}p^{\text{text2image}}(x_{i}\mid c^{\text{text}})\right)q^{\text{VLM}}(x\mid c^{\text{text}})$, as described in Section 4.4. Prompts from Figs. 8 and 6 are taken from GenAI-Bench li2024genai . In Section 4.4 and Fig. 8, we use $T=28$ and image resolution $1024\times 1024$, with $1$ MCMC sampling step per iteration.

### v2

@@ C.1 Flow Models
We require that $\xi$ is monotonic and $\xi(T)=0,\xi(1)=1$. By construction, $p^{(i)}_{T}(x)=\mathcal{N}(0,I)$ is easy to sample from, and $p^{(i)}_{1}(x)=p_{\text{data}}^{(i)}(x)$ encapsulates the prior data distribution of the expert model, fulling our motivation in Section 3.1. For all experiments, we define $\xi$ to be the Euler discretization (Esser et al., 2024; Euler, 1792), with additional alignment to the DDPM scheduler (Ho et al., 2020) for experiments in Section 4.3.
@@
$s_{t}^{(i)}(x)$ $=\frac{-\alpha_{t}v_{t}^{(i)}(x)+\dot{\alpha}_{t}x}{\dot{\sigma_{t}}\sigma_{t}\alpha_{t}-\dot{\alpha}_{t}\sigma^{2}_{t}},$
@@
where $\lambda^{(i)}(x)$ is defined in Section E.1, with all entries of $\sum_{i}\lambda^{(i)}(x)$ to be $1$.
The end-point prediction is available as
$\hat{x}=\frac{\sigma_{t}}{\dot{\alpha}_{t}\sigma_{t}-\dot{\sigma}_{t}\alpha_{t}}v_{t}(x)-\frac{\dot{\sigma}_{t}}{\dot{\alpha}_{t}\sigma_{t}-\dot{\sigma}_{t}\alpha_{t}}x.$ (8)
C.2 Framework Instantiation with Autoregressive Experts
C.3 Approximation for Conditional Sampling
We explain Eq. 6 below. Let $x_{i}$ be any expert. We approximate the distribution $p(x_{i}\mid x_{\text{pa}(i)})$ as
@@ C.3 Approximation for Conditional Sampling
Appendix D Extended Experimental Results
D.1 Graphics-Engine Instructed Image Editing
Fig. 6 contains qualitative samples on the Magic Insert dataset for experiments in Section 4.1.
Figure 6: Comparisons on Character Insertion Task.
D.2 Physical-Simulator-Instructed Video Generation
Qualitative samples are included in Fig. 7.
D.3 Text-to-Image Generation with Regional Composition
Fig. 8 contains qualitative samples on the MIG-Bench dataset for experiments in Section 4.3.
Figure 8: Comparisons on Layout-Conditioned Image Generation Task.
D.4 Ablation on SMC
Method mIoU ($\uparrow$) VQAScore ($\uparrow$) VQAScore-R ($\uparrow$) Wall-Clock Time (min)
[…]
$L=32$ $\mathbf{0.728}$$\pm 0.170$ $\mathbf{0.904}$$\pm 0.092$ $\mathbf{0.881}$$\pm 0.115$ $\sim$25
[…]
Table 4: SMC resampling ablations on the task from Section 4.3. Values in green (red) indicate “No SMC” is better (worse). Results suggest that SMC benefits sampling efficiency. Resampling has relatively marginal computation.
We conduct ablation for SMC with results shown in Table 4.
D.5 Text-to-Image Generation
We further demonstrate the efficacy of distribution composition on the task of text-to-image generation, where we first query an LLM to parse an input text prompt into spatial bounding boxes with regional prompts (Fig. 9 left), aiming to compose the distributions from text-to-image model (FLUX.1 [dev]) conditioned on regional text prompts, one for each region, as generative experts, and from VQAScore (Lin et al., 2024) as a discriminative expert, which gives a scalar value predicting the probability that its input image is aligned with the input text prompt. We compare our methods using $L=16,4$ particles and $1$ particle (which amounts to not using the discriminative expert), a variant using no generative composition (equivalent to best-of-N sampling for vanilla FLUX) while matching the compute budget for 16 particles, and with the vanilla FLUX backbone. In the examples shown in Fig. 9, our compositional approach achieves better text alignment compared to the baselines, and the discriminative expert provides crucial knowledge, e.g., counting, to steer the output distribution.
Figure 9: Text-to-Image Generation.
@@ D.5 Text-to-Image Generation
The parallel sampling process is visualized in Fig. 10. For each scene from Section 4.3, we use an LLM to convert a text prompt into a layout of bounding boxes following the procedure described in Section E.5. The input text prompts and resulting layouts and regional prompts are visualized in Fig. 9. Then, we follow Algorithm 1 and initialize $L=4$ particles and annealed schedule length $T=28$, which are reweighed at three key timesteps $t=21,19,1$ using the discriminative expert, VQAScore (Lin et al., 2024), which measures image-text alignment with a probability score. At $t=21,19$, we binarize the scores with the threshold set to be the median score of the current 4 particles, discard the two below the threshold, and duplicate the remaining two high-score particles two times to maintain 4 particles in each round. At $t=1$, we keep the highest-scoring particle as the final output sample.
@@
Fig. 10 shows end-point predictions to visualize the intermediate samples from the annealing process from $t=28$ to $t=1$. VQAScore measurements (higher means more aligned) are overlayed on the top-left corner for each sample. In the proposed framework, integrating multiple generative experts allows for composing simpler regional prompts, construction of the annealed distributions allows for iterative refinement of samples, and parallel sampling further integrates the knowledge from the discriminative expert and improves output accuracy, as evidenced in the comparisons from Fig. 9.
Appendix E Experiment Details
E.1 Implementation Details
As described in Eq. 4, a generative expert can be defined over the entire scene $x\in\mathbb{R}^{d}$ or a region $x_{i}\in\mathbb{R}^{d^{(i)}}$ with $d^{(i)}\leq d$, in which case the expert has prior over only part of the scene, e.g., a spatial region in Section 4.3. In these cases, we zero-pad the model prediction from $\mathbb{R}^{d^{(i)}}$ to $\mathbb{R}^{d}$.
In other cases, one may restrict the influence of a generative expert to update only a scene region, even when the expert is defined over the full scene $x\in\mathbb{R}^{d}$, e.g., when the expert only has knowledge over foreground object informed by the graphics engine or physics simulators (Sections 4.1 and 4.2). We apply a regional (spatial for images and spatio-temporal for videos) weight $\lambda^{(i)}(x)\in\left[0,1\right]^{d}$ on the velocity prediction $v_{t}^{(i)}(x)$ in these cases, where $\lambda^{(i)}(x)$ is a Gaussian-blurred foreground mask.
E.2 Automatic Evaluation with VLMs
@@
We follow the protocol proposed in PhysGen3D (Chen et al., 2025) to run automatic evaluation for Sections 4.3 and 4.2 using GPT-4o (Hurst et al., 2024). For each scene, all methods’ output images, or uniformly sampled 10 frames when outputs are videos, are sent to GPT-4o together with template prompts as specified in PhysGen3D (Chen et al., 2025). GPT-4o is asked to give scores from 0 to $1$ for each video on all metrics. The scores for the same metric are averaged across all scenes.
E.3 Graphics-Engine-Instructed Image Editing
@@
In Section 4.1, the product distribution is defined as $p(x)\propto p^{\text{depth2image}}(x\mid c^{\text{text}},c^{\text{depth}})p^{\text{imagefill}}(x\mid c^{\text{text}},c^{\text{image}},c^{\text{mask}})$. We use $T=28,L=1$, with 1 MCMC sampling step and 2 conditional sampling updates (Eq. 6) per iteration. The spatial resolution is $960\times 1664$ for the graphics engine dataset and $1024\times 1024$ for Magic Insert dataset.
For the graphics engine dataset, assets are rendered in Genesis [^2] https://github.com/Genesis-Embodied-AI/Genesis.
For Magic Insert data preprocessing, we use SAM2 (Ravi et al., 2024) to obtain the segmentation masks of foreground characters, and feed the images of segmented characters, resized to the bottom-left quadrant and overlaid with backgrounds, into GPT o4-mini for image captions to serve as text prompts for evaluated methods. Magic Insert results are downloaded from the official demo page. Add-it receives background images from the dataset as inputs, together with the GPT-generated captions (of overlaid images) appended with “{character} is at the bottom left quarter” to indicate the desired locations. SDEdit takes in input overlaid images and adds Gaussian noise with standard deviation $0.5$ before denoising.
E.4 Physical-Simulator-Instructed Video Generation
@@
In Section 4.2, for flow models, the product distribution is defined as $p(x)\propto p^{\text{depth2video}}(x\mid c^{\text{text}},c^{\text{depth}})p^{\text{image2video}}(x\mid c^{\text{text}},c^{\text{image}})$. We use $T=30,L=1$ with 1 MCMC sampling step per iteration. We use no conditional sampling updates for this task due to GPU memory constraint. Each generated video has $81$ frames and spatial resolution $480\times 832$.
For autoregressive models, the product distribution is defined as $p(x)\propto p^{\text{sim}}(x\mid c^{\text{sim}})p^{\text{image2video}}(x\mid c^{\text{text}},c^{\text{image}})$ with $p^{\text{sim}}(x\mid c^{\text{sim}})$ defined in Section 4.2. We approximate sampling from this distribution with first sampling from $p^{\text{image2video}}$, then do 8 gradient updates with respect to $\mathcal{L}_{2}$ loss $\|x-c^{\text{simulation}}\|^{2}$. Results in Fig. 5 use $81$ frames per sequence and spatial resolution $512\times 768$. We set $L=1$, with 1 MCMC sampling step per iteration for this experiment.
E.5 Text-to-Image Generation with Regional Control
For Section 4.3, the product distribution is $p(x)\propto\left(\prod_{i}p^{\text{depth2image}}(x_{i}\mid c^{\text{text}},c^{\text{depth}})\right)q^{\text{VLM}}(x\mid c^{\text{text}})$. Here, $p^{\text{depth2image}}$ is a regional generative experts (FLUX-Depth), which predict scores for bounding-box-cropped images conditioning on regional text prompts and on regional depth maps cropped from global depth maps predicted using the first stage of 3DIS-FLUX (Zhou et al., 2025) (layout-to-depth generation); $q^{\text{VLM}}$ a discriminative expert VQAScore, where for each image sample, it assigns the summed score computed over a full image and its regions cropped with input bounding boxes as the reward. We use $T=28$ and image resolution $1024\times 1024$, with $1$ MCMC sampling step and $2$ conditional sampling steps per iteration.